You are an entrepreneur starting a biotechnology firm. If your research is​ successful, the technology can be sold for $33 million. If your research is​ unsuccessful, it will be worth nothing. To fund your​ research, you need to raise $2.3 million. Investors are willing to provide you with $2.3 million in initial capital in exchange for 25% of the unlevered equity in the firm.

a. What is the total market value of the firm without​ leverage?

b. Suppose you borrow $0.7 million. According to​ MM, what fraction of the​ firm's equity will you need to sell to raise the additional $1.6 million you​ need?

c. What is the value of your share of the​ firm's equity in cases ​(a​) and ​(b​)?

Case ​(a​) is ​$ ____?

Case (b) is $ ___?
Answer #1

Ans:

(a) Total market value of the firm without leverage = $ 2.3 million / 25 %

= $ 2.3 * 4

= $ 9.2 million

(b) MM says total value of firm is still $ 9.2 million . 0.7 million of debt implies total value of equity is ( 9.2 - 0.7) = 8.5 million .

0.7 / 8.5 million = 8.235 %

(c) 25% of 9.2 million = 2.3 million

(100 - 8.235 % ) * 8.5 million = 7.8 million .